Given that O(0,0) P(x,4) and Q(8,1) are three points. If OP=OQ Find possible values of x

Grade:10

1 Answers

Arun
5 years ago
Dear student
 
If OP = OQ 
Then
Sqrt( x² +16) = sqrt (64 + 1)
x² +16 = 64 +1
x² = 49
x = ± 7
Regards
Arun (askIITians forum expert)